Assume that P(E) = 0.15 and P(F) = 0.48. If E and F are independent, find P(E and F).
Answer


Answer by stanbon(75887)

Assume that P(E) = 0.15 and P(F) = 0.48. If E and F are independent, find P(E and F).
---
P(E AND F) = P(E)*P(F) = 0.15*0.48 = 0.0720
==============
Cheers,
Stan H.
==============